OpenStudy (anonymous):

can any buddy help ill give u meddals and fan ok 5 questions

OpenStudy (anonymous):

Which is a solution to the equation y = 2x – 5? A. (2, 1) B. (5, –5) C. (3, 11) D. (1, –3)

OpenStudy (anonymous):

You have to plug in the x and y values. y = 2x - 5 (1, -3) is the only one that works. -3=2(1)-5 -3=2-5 -3=-3 11=2(3)-5 11=6-5 11 does not equal 1. 1=2(2)-5 1=4-5 1 does not equal -1. -5=2(5)-5 -5=10-5 -5 does not equal 5.

OpenStudy (anonymous):

so its d

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok next question

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

 A. (–4, –7) B. (3, 14) C. (2, –1) D. (1, 8)

OpenStudy (anonymous):

i think it is c

OpenStudy (anonymous):

C sounds the most logical.

OpenStudy (anonymous):

so next question

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

What is the value of y so that the ordered pair (5, y) is a solution to the equation –5x + y = –40? A. –65 B. 9 C. –15 D. 7

OpenStudy (anonymous):

i think it is a

OpenStudy (anonymous):

What is the value of y so that the ordered pair (5, y) is a solution to the equation -5x + y = -40? Answer: C. -15

OpenStudy (anonymous):

Do you understand how I did it?

OpenStudy (anonymous):

kinda not allot tho

OpenStudy (anonymous):